OpenStudy (anonymous):

is this table an exponential function? x = 1, 2, 3, 4 y = -1, -8, -27, -64

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yes is the answer

OpenStudy (unanimoose):

Actually it is not. An exponential function will have y values that look something like this: 1, 4, 16, 64

OpenStudy (unanimoose):

Notice this: 1 * 4 = 4 4 * 4 = 16 16 * 4 = 64

OpenStudy (unanimoose):

Now when we compare that with the given y value here: -1 * 8 = -8 -8 * 8 = -64 As you can see, -8 * 8 does NOT = -27.

OpenStudy (unanimoose):

So the values are not consistent making it not exponential.

OpenStudy (anonymous):

oh okay, i get it thank you very much !!!

OpenStudy (unanimoose):

Np.